# Denoising Neural Reranker for Recommender Systems

### v2

@@ Abstract
For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then the recommender calls a slower but more sophisticated reranking model that refines the item list exposure to the user. To consistently optimize the two-stage retrieval reranking framework, most efforts have focused on learning reranker-aware retrievers. In contrast, there has been limited work on how to achieve a retriever-aware reranker. In this work, we provide evidence that the retriever scores from the previous stage are informative signals that have been underexplored. Specifically, we first empirically show that the reranking task under the two-stage framework is naturally a noise reduction problem on the retriever scores, and theoretically show the limitations of naive utilization techniques of the retriever scores. Following this notion, we derive an adversarial framework DNR that associates the denoising reranker with a carefully designed noise generation module. The resulting DNR solution extends the conventional score error minimization loss with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones. We conduct extensive experiments on three public datasets and an industrial recommender system, together with analytical support, to validate the effectiveness of the proposed DNR.
@@ 1 Introduction
Recommender systems aim to personalize a list of items that maximizes user engagement signals, such as purchases in e-commerce, video plays in video platforms, and link sharing in social networks. To achieve a feasible solution with large-scale candidate pools and user requests, industrial solutions Covington et al. (2016); Ricci et al. (2022); Liu et al. (2025) typically employ a multi-stage architecture comprising a retriever and a reranker. The retriever, which may further decompose into several more sub-stages of retrieval and ranking modules, efficiently filters a large item pool down to a manageable candidate set of relevant items. Then the reranker adopts a more sophisticated but more accurate model that refines the candidate item set into an optimal ordering for the list-wise exposure. To optimize the two-stage framework towards a consistent goal of recommendation (e.g., aligning true user behavior or preference) across all stages, the majority of efforts have been focusing on learning a reranker-aware retriever Gallagher et al. (2019); Xu et al. (2024a). However, there is limited work on how to achieve a retriever-aware reranker.
@@
Formally, the goal of the latter reranking task Carbonell & Goldstein (1998); Jin et al. (2008); Pei et al. (2019); Liu et al. (2022) is to learn an ordering function over the retrieved candidates that maximizes user responses on the top-reranked items. Early studies propose to re-score the initial retrieval items Jin et al. (2008); Ai et al. (2018); Pei et al. (2019) by modeling the mutual influences among items in the candidate list. Recent state-of-the-art approaches find it more effective to formulate the reranking task as a list generation problem Shi et al. (2023); Ren et al. (2024); Lin et al. (2024), where the lists with better list-wise rewards are given higher generation probability. Yet, existing methods largely overlook the potential of initial retriever scores in the input space, which may offer rich prior information to enhance the reranking performance. In practice, a straightforward approach to leverage initial retriever scores is to include them as input features directly, which shows promise to improve reranking performance as in Figure 1(a). Still, as we will illustrate in section 3.2, this naive solution might be far from exploiting this retriever information effectively to align with user feedback, which is addressed by the following contributions in this work.
@@
Formulating reranking as a noise reduction problem: The retriever stage typically employs a simpler model than the reranker to process large candidate pools, given computational budget constraints. In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(a), 1(b)). What follows is an empirical assumption that there is higher noise in the retriever scores than the reranker scores, which naturally indicates a noise reduction process of the reranking stage. This noise reduction pattern is more evident when we observe the changes in the embedding representations of ground truth target items. We show evidence by comparing Figure 1(c) and Figure 1(d), where the same set of retrieved candidates is not distinguished well by the retriever, but the target items are better predicted by the reranker, in terms of the embedding space. This further proves that the reranker behaves like a denoising model on the retriever’s noise. However, directly aligning the user behavior on the condition of the observed retriever score, ignoring the uncertainty of this noisy signal, may cause a misalignment with the user behavior alignment goal in terms of the entire system. Thus, we argue that learning a noise-aware reranker is crucial for achieving a retriever-aware reranker.
@@
(c) Target item and selected candidates embedding in the retriever stage.
(d) Noise reduction process in the item embedding perspective.
Figure 1: (a) The informative retriever score and (b-d) noise reduction nature of reranking. The reranker in this example uses transformer model to select top-20 items from the candidates. “Rerank w/ s” represents using retriever scores as additional item features for reranker. The blue shaded circles in (c) and (d) represent the retriever’s selection of candidate items. The green shaded circle in (d) represents the reranker’s selection of the items of exposure.
@@ 2.1 Multi-stage Recommendation
Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems Qin et al. (2022); Zheng et al. (2024); Liu et al. (2022) (also known as cascade ranking Liu et al. (2017); Gallagher et al. (2019) process) to balance accuracy and efficiency under strict latency constraints. Each stage progressively narrows down the candidate set while employing increasingly sophisticated models. While most existing work focuses on independently optimizing individual stages (e.g., retrieval Wang et al. (2011; 2017) or reranking stage Liu et al. (2022); Pei et al. (2019)), recent studies explore joint optimization across the entire pipeline Gallagher et al. (2019); Qin et al. (2022); Xu et al. (2024b). These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance Zhao et al. (2023); Xu et al. (2024a). However, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
@@ 2.2 Reranking Strategies
In sequential recommendation, re-ranking Pei et al. (2019); Huzhang et al. (2023); Ai et al. (2018) aims to refine the initial ranking list generated by a previous retrieval stage. Existing re-ranking approaches can be broadly categorized into the following paradigms. Single-point re-scoring methods Kang & McAuley (2018); Zhuang et al. (2018) independently predict a refined score for each item and re-rank them accordingly. While efficient, they ignore dependencies between items, potentially leading to suboptimal list-level performance. Unlike single-point approaches, list-refinement methods Pei et al. (2019); Ai et al. (2018); Xi et al. (2022) explicitly model mutual influences between items by treating the pre-ranked candidate list as input. Techniques such as Transformer-based architectures Pei et al. (2019) capture user preference to optimize the list holistically. Generator-evaluator methods Huzhang et al. (2023); Shi et al. (2023) first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one. They can effective generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop. Emerging techniques leverage diffusion models Lin et al. (2024); Li et al. (2024) to generate refined item lists by iteratively denoising a ranking distribution, which excel in capturing complex user preferences.
@@ 2.3 Adversarial Learning
Adversarial learning Lowd & Meek (2005) has emerged as a powerful technique for data augmentation in recommendation systems, improving the models robustness and generalization. By leveraging generative adversarial networks (GANs) Wang et al. (2017) or adversarial training He et al. (2018); Tang et al. (2020), these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness. For instance, IRGAN Wang et al. (2017) employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real vs. synthetic interactions) to improve ranking performance. Similarly, AdvIR Park & Chang (2019) introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations. These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness. Our method is designed to add noise by adversarial learning, which can enhance the denoising models (rerankers)’ effectiveness in denosing the pre-ranking scores to align with user feedback.
@@ 3.2 Analytical Limitation of Direct Optimization Methods
where $p_{\text{z|x}}$ represents the ground-truth user feedback probability and $p_{\text{x|z}}$ represents the posterior distribution of retriever scores conditioned on user feedback, both terms are not related to the reranker $q_{\theta}$ and they are merely determined by the retriever prior and the user. Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1). However, these two terms are not fully controlled by the direct optimization framework, potentially causing significant misalignment between Eq.(2) and Eq.(1). This will lead to a discrepancy between the final reranking results and user feedback, and thus, the suboptimal recommendation performance.
@@ 3.3 Implementing Noise Generator for the Posterior
To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a noise generator $f_{\phi}(\cdot|\mathbf{z}_{u})$. Formally, we generate noise with $\bm{\epsilon}\sim f_{\phi}$ and add it to the user feedback to form the synthetic retriever score posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$, taking insight from the reparameterization trick Kingma & Welling (2014):
@@
  • Heuristic Generator: This option uses a non-parametric noise generator $f_{\phi}^{\text{heuristic}}$ with a manually designed distribution. For example, we can adopt random Gaussian noise similar to Ho et al. (2020); Li et al. (2024), but we notice that this is not a good representation of the noise in retriever scores. Intuitively, given the binary user feedback signals, a natural assumption of the retriever prior is the Beta distribution, where we can derive the heuristic solution as $\bm{\epsilon}\sim\text{Beta}(\alpha,\beta)$. For simplicity, one may set $\alpha=\beta=0.5$ to approximate the uncertainty. In practice, we find that it is also beneficial to set $\alpha$ and $\beta$ that align with the observed noise distribution in the data.
  • Model-based Generator: While the heuristic generator is effective in our empirical study, we are skeptical of whether the heuristic noise generator is optimal for augmenting the retriever score distribution. To achieve a more precise estimation, we design a learnable model $f_{\phi}^{\text{model}}$, with parameter $\phi$, that will be optimized along with the reranker (details in the next section). The generated noise $\bm{\epsilon}$ should also be aligned with the observed distribution, and it also participates in an adversarial training objective in order to better explore the retriever score space, which in turn improves the reranker’s effectiveness. As we will show in Section 4.3, this choice is empirically better than the heuristic generator.
@@ 3.4 Overall Learning Framework
In this section, we show how to optimize the user feedback alignment goal i.e., Eq.(1) by simultaneously learning the denoising reranker and the adversarial noise generator. Specifically, we can decompose the negative log-likelihood objective into three loss terms (derivation in Appendix B.1):
@@
We implement the model-based generator $f_{\phi}^{\text{model}}$ by a 2-layer MLP and directly output $\mathbf{x}_{u}^{\prime}$ using Eq.(4), so that the end-to-end learning becomes feasible for Eq.(7). In cases where the noise generator is not directly optimizable, one may have to consider other alternatives like reparameterization tricks Kingma & Welling (2014) or policy gradient Silver et al. (2014). Notably, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker. In such cases, the synthetic scores are unrealistically different from the real retriever scores, diverging from the user feedback, inducing small $p_{\text{z|x}}$ and large $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
@@ 4 Experiments
RQ1: How does DNR perform in the reranking tasks of recommendation systems compared with leading rerankers? RQ2: How well is the denoising formulation compared to alternative methods of leveraging retriever scores? RQ3: Can the denoising formulation generalize to other reranker backbones, rather than PRM? RQ4: What is the performance difference between the heuristic noise generator and the adversarial model-based generator? RQ5: How does different objectives in DNR contribute to the performance? RQ6: How sensitive is DNR to the hyperparameters $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$?
@@ 4.1 Expermental Settings
Datasets. We conduct experiments on three real-world datasets, including ML-1M Harper & Konstan (2016), Kuaivideo Liu et al. (2025), and Amazon-books Ni et al. (2019). The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix D.1. For the online experimental verification, please refer to Appendix E.6.
Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec Kang & McAuley (2018), Caser Tang & Wang (2018), GRU4Rec Hidasi et al. (2016), and MiDNN Zhuang et al. (2018)), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM Ai et al. (2018), SetRank Pang et al. (2020), PRM Pei et al. (2019), and MIR Xi et al. (2022)), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank Huzhang et al. (2023), Pier Shi et al. (2023), and NAR4Rec Ren et al. (2024)), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec Li et al. (2024) and DCDR Lin et al. (2024)) which utilize diffusion models to generate items or lists for recommendation. Detailed information for our baselines is presented in Appendix D.3.

### v3

@@ Abstract
For multi-stage recommenders in industry, a user request would first trigger a simple and efficient retriever module that selects and ranks a list of relevant items, then the recommender calls a slower but more sophisticated reranking model that refines the item list exposure to the user. To consistently optimize the two-stage retrieval reranking framework, most efforts have focused on learning reranker-aware retrievers. In contrast, there has been limited work on how to achieve a retriever-aware reranker. In this work, we provide evidence that the retriever scores from the previous stage are informative signals that have been underexplored. Specifically, we first empirically show that the reranking task under the two-stage framework is naturally a noise reduction problem on the retriever scores, and theoretically show the limitations of naive utilization techniques of the retriever scores. Following this notion, we derive an adversarial framework DNR that associates the denoising reranker with a carefully designed noise generation module. The resulting DNR solution extends the conventional score error minimization loss with three augmented objectives, including: 1) a denoising objective that aims to denoise the noisy retriever scores to align with the user feedback; 2) an adversarial retriever score generation objective that improves the exploration in the retriever score space; and 3) a distribution regularization term that aims to align the distribution of generated noisy retriever scores with the real ones. We conduct extensive experiments on three public datasets and an industrial recommender system, together with analytical support, to validate the effectiveness of the proposed DNR. The code is released at https://github.com/maowenyu-11/DNR.
@@ 1 Introduction
Recommender systems aim to generate personalized item lists and maximize users’ engagement. For large-scale item pools, industrial solutions Ricci et al. (2022); Liu et al. (2025) typically use a two-stage retriever-reranker architecture. The retriever efficiently narrows down a large item pool to a manageable relevant candidate set. The reranker then uses a more sophisticated model to reorder the candidates into an optimal list-wise order. To optimize the two-stage framework towards a consistent goal of recommendation (e.g., aligning with user behavior or preference) across all stages, the majority of efforts have been focusing on learning a reranker-aware retriever Gallagher et al. (2019); Xu et al. (2024a). However, there is limited work on how to achieve a retriever-aware reranker.
@@
For the latter reranker Carbonell and Goldstein (1998); Jin et al. (2008); Pei et al. (2019); Liu et al. (2022), early studies propose to re-score the retrieval items Jin et al. (2008); Ai et al. (2018); Pei et al. (2019) by modeling the mutual influences among candidate items. Recent state-of-the-art approaches find it more effective to formulate the reranking task as a list generation problem Shi et al. (2023); Ren et al. (2024); Lin et al. (2024), where the lists with better list-wise rewards are given higher generation probability. Yet, existing methods largely overlook the potential of initial retriever scores, which may offer rich prior information for the reranker stage. In practice, a straightforward approach to leverage initial retriever scores is to directly include them as extra input features of the model, which shows promise to improve reranking performance as in Figure 1(a). Still, as we will illustrate in section 3.2, this naive solution might be far from exploiting this retriever information effectively to align with user feedback, which is addressed by the following contributions in this work.
@@
Formulating reranking as a noise reduction problem: The retriever stage typically employs a simpler model than the reranker to process large candidate pools, given computational budget constraints. In contrast, the reranker, operating on a much smaller candidate set, can use a more sophisticated architecture and thus achieves significantly higher accuracy (Figure 1(a), 1(b)). What follows is an empirical assumption that there is higher noise in the retriever scores than the reranker scores, which naturally indicates a noise reduction process of the reranking stage. This noise reduction pattern is also evident when we observe the changes in the embedding representations of ground truth target items. We show evidence by comparing Figure 1(c) and Figure 1(d), where the same set of retrieved candidates is not distinguished well by the retriever, but the target items are better predicted by the reranker, in terms of the embedding space. The reranker’s noise is more concentrated near zero than the retriever’s (Figure 1(e)), further confirming the reranker acts as a denoising model for the retriever’s noise. However, due to the uncertainty inherent in noisy retriever scores, denoising based on these scores may misalign with the system-wide goal of aligning with user behavior. Thus, we argue that learning a noise-aware reranker is crucial for achieving a retriever-aware reranker.
@@
(c) Target item and selected candidates embedding in retriever.
(d) Noise reduction process in item embedding perspective.
(e) Noise density distribution of the reranker and retriever.
Figure 1: (a) The informative retriever score and (b-d) noise reduction nature of reranking. The reranker in this example uses transformer model to select the top-20 items from the candidates. “Rerank w/ s” represents using retriever scores as additional item features for the reranker. The black shaded circles in (c) and (d) represent the retriever’s selection of candidate items, and the green shaded circle in (d) represents those of the reranker for exposure. (e) compares the noise distribution (distance between predicted scores and ground-truth labels) of the retriever and the reranker.
@@ 2.1 Multi-stage Recommendation
Modern content-intensive web services with large candidate pools typically employ multi-stage recommender systems Qin et al. (2022); Zheng et al. (2024); Liu et al. (2022) (also known as cascade ranking Liu et al. (2017); Gallagher et al. (2019) process) to balance accuracy and efficiency under strict latency constraints. Each stage progressively narrows down the candidate set while employing increasingly sophisticated models. While most existing work focuses on independently optimizing individual stages (e.g., retrieval Wang et al. (2011; 2017) or the reranking stage Liu et al. (2022); Pei et al. (2019)), recent studies explore joint optimization across the entire pipeline Gallagher et al. (2019); Qin et al. (2022); Xu et al. (2024b). These approaches leverage later-stage models to guide earlier-stage learning, improving overall coherence and performance Zhao et al. (2023); Xu et al. (2024a). In contrast, our work takes a complementary direction—–using earlier-stage scores to regulate later-stage optimization.
@@ 2.2 Reranking Strategies
In sequential recommendation, re-ranking Pei et al. (2019); Huzhang et al. (2023); Ai et al. (2018) aims to refine the initial ranking list generated by a previous retrieval stage. Existing re-ranking approaches can be broadly categorized into the following paradigms. Single-point re-scoring methods Kang and McAuley (2018); Zhuang et al. (2018) independently predict a refined score for each item and re-rank them accordingly. While efficient, they ignore dependencies between items, potentially leading to suboptimal list-level performance. Unlike single-point approaches, list-refinement methods Pei et al. (2019); Ai et al. (2018); Xi et al. (2022) explicitly model mutual influences between items by treating the pre-ranked candidate list as input. Techniques such as Transformer-based architectures Pei et al. (2019) capture user preferences to optimize the list holistically. Generator-evaluator methods Huzhang et al. (2023); Shi et al. (2023) first generate multiple candidate lists and then evaluate them using a learned utility function to select the best one. They can effectively generate a high-quality recommendation list but suffer from the increasing computational cost due to the generation-evaluation loop. Emerging techniques leverage diffusion models Mao et al. (2025c; a; d) to generate refined item lists by iteratively denoising a ranking distribution, which excels in capturing complex user preferences.
@@ 2.3 Adversarial Learning
Adversarial learning Lowd and Meek (2005) has emerged as a powerful technique for data augmentation in recommendation systems, improving the model robustness and generalization. By leveraging generative adversarial networks (GANs) Wang et al. (2017) or adversarial training He et al. (2018); Tang et al. (2020), these methods synthesize high-quality user-item interactions or perturb existing data to enhance model robustness. For instance, IRGAN Wang et al. (2017) employs a minimax game between a generator (sampling hard negative items) and a discriminator (distinguishing real from synthetic interactions) to improve ranking performance. Similarly, AdvIR Park and Chang (2019) introduces adversarial perturbations to embedding spaces, forcing the model to learn more generalizable representations. These methods demonstrate that adversarial learning can not only augment training data but also enhance models’ robustness. Our method is designed to add noise by adversarial learning, which can enhance the reranker’s effectiveness in denoising the pre-ranking scores to align with user feedback.
@@ 3.2 Analytical Limitation of Direct Optimization Methods
where $p_{\text{z|x}}$ represents the ground-truth user feedback probability and $p_{\text{x|z}}$ represents the posterior distribution of retriever scores conditioned on user feedback, both terms are not related to the reranker $q_{\theta}$, and they are merely determined by the retriever prior and the user. Intuitively, smaller values of $L_{1}$ and $L_{2}$ make the solution of Eq.(2) more aligned with the goal of Eq.(1). However, these two terms are not fully controlled by the direct optimization framework, potentially causing significant misalignment between Eq.(2) and Eq.(1). This will lead to a discrepancy between the final reranking results and user feedback, and thus, the suboptimal recommendation performance.
@@ 3.3 Implementing Noise Generator for the Posterior
To overcome the aforementioned limitations, we propose to complement the denoising reranker $q_{\theta}$ with a noise generator $f_{\phi}(\cdot|\mathbf{z}_{u})$. Formally, we generate noise with $\bm{\epsilon}\sim f_{\phi}$ and add it to the user feedback to form the synthetic retriever score posterior $p_{\phi}(\mathbf{x}|\mathbf{z}_{u})$, taking insight from the reparameterization trick Kingma and Welling (2014):
@@
  • Heuristic Generator: This option uses a non-parametric noise generator $f_{\phi}^{\text{heuristic}}$ with a manually designed distribution. For example, we can adopt random Gaussian noise similar to Ho et al. (2020); Li et al. (2024), but we notice that this is not a good representation of the noise in retriever scores. Intuitively, given the binary user feedback signals, a natural assumption of the retriever prior is the Beta distribution, i.e., , the conjugate prior and the posterior for binary variables (i.e., bernoulli or binomial distribution) are both beta distribution, where we can derive the heuristic solution as $\bm{\epsilon}\sim\text{Beta}(\alpha,\beta)$. For simplicity, one may set $\alpha=\beta=0.5$ to approximate the uncertainty. In practice, we find that it is also beneficial to set $\alpha$ and $\beta$ that align with the observed noise distribution in the data.
  • Model-based Generator: While the heuristic generator is effective in our empirical study, we are skeptical of whether the manually designed heuristic noise generator is optimal for augmenting the retriever score distribution. Since the prior noise distribution is uncertain, it is necessary to learn this distribution adaptively and personally. To achieve a more precise estimation, we design a learnable model $f_{\phi}^{\text{model}}$, with parameter $\phi$, that will be optimized along with the reranker (details in the next section). The generated noise $\bm{\epsilon}$ should also be aligned with the observed distribution, and it also participates in an adversarial training objective in order to better explore the retriever score space, which in turn improves the reranker’s effectiveness. As the noise is learned, $\lambda_{c}$ requires no manual adjustment and can be treated as 1. As we will show in Section 4.3, this choice is empirically better than the heuristic generator.
@@ 3.4 Overall Learning Framework
In this section, we show how to optimize the user feedback alignment goal, i.e., Eq.(1) by simultaneously learning the denoising reranker and the adversarial noise generator. Specifically, we can decompose the negative log-likelihood objective into three loss terms (derivation in Appendix B.1):
@@
We implement the model-based generator $f_{\phi}^{\text{model}}$ by a 2-layer MLP and directly output $\mathbf{x}_{u}^{\prime}$ using Eq.(4), so that the end-to-end learning becomes feasible for Eq.(7). In cases where the noise generator is not directly optimizable, one may have to consider other alternatives like reparameterization tricks Kingma and Welling (2014) or policy gradient Silver et al. (2014). Notably, $\mathcal{L}_{\text{adv}}$ does not encourage over-exploration of $\mathbf{x}_{u}^{\prime}$, which may negatively impact the learning of the reranker. In such cases, the synthetic scores are unrealistically different from the real retriever scores, diverging from the user feedback, inducing small $p_{\text{z|x}}$ and large $q_{\theta}(\mathbf{z}_{u}|\mathbf{x})/p_{\text{z|x}}(\mathbf{z}_{u}|\mathbf{x})$, violating $\mathcal{L}_{\text{adv}}$.
@@ 4 Experiments
  • RQ1: How does DNR perform compared with leading rerankers in recommendation systems?
  • RQ2: How does the denoising formulation compare to other methods of using retriever scores?
  • RQ3: Can the denoising formulation generalize to reranker backbones other than PRM?
  • RQ4: How does the heuristic noise generator’s performance compare to that of the adversarial model-based generator?
  • RQ5: How does different objectives in DNR contribute to the performance?
  • RQ6: How sensitive is DNR to the hyperparameters $\lambda_{c}$, $\lambda_{m}$, and $\lambda_{e}$?
@@ 4.1 Expermental Settings
Datasets. We conduct experiments on three real-world datasets, including ML-1M Harper and Konstan (2016), Kuaivideo Liu et al. (2025), and Amazon-books Ni et al. (2019). The detailed description and the statistics of the processed datasets for both the retriever and the reranker stages are presented in Appendix D.1. For the online experimental verification, please refer to Appendix E.6.
Baselines: We compare the performance of DNR with four categories of recommendation methods in our reranking task, including traditional recommenders (i.e., SASRec Kang and McAuley (2018), Caser Tang and Wang (2018), GRU4Rec Hidasi et al. (2016), and MiDNN Zhuang et al. (2018)), which individually predict the scores of candidate items and rank them accordingly; list-refinement methods (i.e., DLCM Ai et al. (2018), SetRank Pang et al. (2020), PRM Pei et al. (2019), and MIR Xi et al. (2022)), which encode the mutual inference in candidate items and refine the rankings based on this information; generator-evaluator methods (i.e., EGRerank Huzhang et al. (2023), Pier Shi et al. (2023), and NAR4Rec Ren et al. (2024)), which generate multiple reranked item lists and evaluate them to select the best one for users; and diffusion-Based methods (i.e., DiffuRec Li et al. (2024) and DCDR Lin et al. (2024)) which utilize diffusion models to generate items or lists for recommendation. Detailed information for our baselines is presented in Appendix D.3.